Pronghorn Company reported the following information for November and December 2019. November December Cost of goods purchased $506,000 $ 690,000 Inventory, beginning-of-month 105,000 92,600 Inventory, end-of-month 92,600 ? Sales revenue 810,000 964,000 Pronghorn’s ending inventory at December 31 was destroyed in a fire.

Compute the gross profit rate for November.

Using the gross profit rate for November, determine the estimated cost of inventory lost in the fire.
Answer #1

PART1- Gross Profit Rate for November 2019:

                    Gross Profit Rate= Gross Profit/Sales*100

                    Cost of Goods Sold= Opening Inventory + Purchase – Closing Inventory

                                                     = $105,000 + 506,000 – 92,600

                                                     = $518,400

                   Gross Profit = Sales – COGS = $810,000 - $518,400 = $291,600

                  Gross Profit Ratio= 291,600/810,000*100 = 36%

PART2- First we need to find the COGS for December 2019. Using the same Gross profit Ratio i.e. 36%, Gross Profit For the December 2019 is 36% of 964,000 = $347,040 and

                                     COGS = Sales – Profit = 964,000 – 347,040 = $616,960

                  Cost of Goods Sold= Opening Inventory + Purchase – Closing Inventory

                                      616,960 = 92,600 + 690,000 – Closing Inventory
                                            Closing Inventory = $165,640                           

The Inventory Lost By fire in Dec 2019 $165,640